### v1

@@ 5 Ground State Optimization
In quantum information science an important task relates to finding the eigenstate (‘ground state’) and the smallest eigenvalue (‘ground-state energy’) of a Hermitian operator $A$ that describes the energy of the system (i.e. the Hamiltonian). For this setting, we leave the general notation $f:M\to\mathbb{R}$ above and specialise to the cost function over the unitary group (or orbit) $J:U\to\mathbb{R}$ we aim to minimize. For finding the smallest expectation value of $A$ w.r.t. the state $\rho$, the cost function is given by
$J(U):=\operatorname{tr}(AU\rho U^{*})=\langle A,\operatorname{Ad}_{U}(\rho)\rangle,$ (11)
@@
where $\rho\in\mathfrak{pos}_{1}(n)$ is the initial state of the system and $U$ is a unitary transformation that describes a quantum circuit (writing $U^{*}$ for the complex conjugate transpose).
@@
By choosing an appropriate eigenbasis of $A$ we may assume that $A$ is diagonal with eigenvalues in non-increasing order. The function we want to optimize is of the form of Eqn. (11). Many properties of this function can be found in [22] in the more general setting of semisimple Lie groups. We recall the relevant properties here, adapted to our setting:
@@
Assume that $\rho$ and $A$ are diagonal  [^6] This amounts to a shift in the function $f$ and an appropriate choice of basis.. Then the following holds.
@@
    $U$ is a critical point of $f$ if and only if $[A,\operatorname{Ad}_{U}(\rho)]=0$. Hence the critical set of $f$ is equal to
    $\mathsf{SU}(n)_{A}\,\mathcal{S}_{n}\,\mathsf{SU}(n)_{\rho},$
    where $\mathcal{S}_{n}$ is the set of $n$ by $n$ permutation matrices. In particular it is a disjoint union of finitely many compact connected submanifolds.
@@
    The Hessian at a critical point $U$ is given by
@@
    and in particular $f$ is Morse–Bott. [^7] Note that the function $f$ is never Morse, since the maximal torus in $\mathsf{SU}(n)$ stabilizes $A$ and $\rho$.
@@
i: See Lemma 1.1 and Propositions 1.2 and 1.3 of [22]. ii: See Proposition 1.4 and Corollary 1.5 of [22]. iii: See Remark 1.6 of [22]. ∎
Definition 2.
[26][Thm. 3] A unitary representation $\pi:G\to U(n)$, $g\mapsto\pi(g)=:U_{g}$ a finite group $G$ is called a unitary $t$-design with $t\in\mathbb{N}_{0}$ if one of the following equivalent conditions is satisfied:
@@
    where $\mathbb{C}_{t}[Z,\bar{Z}]$ denotes the set of all polynomials in $Z:=(z_{ij})_{i,j=1,\dots n}$ and $\bar{Z}:=(\bar{z}_{ij})_{i,j=1,\dots n}$, which are $t$ homogeneous in $Z$ and $\bar{Z}$., i.e. [^8] This is equivalent to $p(\lambda Z,\bar{\lambda}\bar{Z})=|\lambda|^{2t}p(Z,\bar{Z})$. with $p(\lambda Z,\mu\bar{Z})=\lambda^{t}\mu^{t}p(Z,\bar{Z})$.
@@
    For all $H\in\mathbb{C}^{n^{t}\times n^{t}}$ one has the equality
@@
    The tensor square representation $\pi\otimes\pi:G\to U(n)\otimes U(n)$, $g\mapsto U_{g}\otimes U_{g}$ acting on $\mathbb{C}^{n}\otimes\mathbb{C}^{n}$ has exactly two irreducible components, namely $\mathrm{Sym}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ and $\mathrm{Alt}(\mathbb{C}^{n}\otimes\mathbb{C}^{n}):=\mathbb{C}^{n}\wedge\mathbb{C}^{n}$.
@@
A more general notion of $t$-designs [20, 21] (see also [26]) allows any finite subset of $U(n)$ which satisfies (a), (b) or, equivalently,(c). However, since almost all known examples are constructed via group representations, we focus here on this restricted approach which is sometimes called group design. Moreover, the reader should note the following to facts: (i) One can assume without loss of generality that $\pi$ is faithful (i.e., one-to-one) because it is straightforward to show that $[g]\mapsto\pi(g)$, $[g]\in G/\mathrm{ker}\pi$ yields a $t$-design whenever $\pi$ is a $t$-design. (ii) Every $t$-design is also $t^{\prime}$-design for $t^{\prime}\leq t$. This follows readily from condition (b) by choosing $H$ of the form $H^{\prime}\otimes I_{n}\otimes\cdots\otimes I_{n}$.
@@
Let $\pi:G\to U(n)$ be any unitary representation. Then $\pi$ yields a $2$-design in the sense of Def. 2 if and only if $\pi$ acts irreducibly on $i\mathfrak{su}(n)$ via conjugation, i.e. the representation $\hat{\pi}:G\to U\big{(}i\mathfrak{su}(n)\big{)}$, $\pi(g)\,H:=U_{g}HU_{g}^{*}$ is irreducible.
@@
First, $\hat{\pi}$ obviously “extends” to $\tilde{\pi}:G\to U\big{(}\mathbb{C}^{n\times n}\big{)}$, $\tilde{\pi}(g)\,A:=U_{g}AU_{g}^{-1}$. Then a straightforward computation shows that the representations $\tilde{\pi}$ and $g\mapsto U_{g}\otimes\bar{U}_{g}$ are equivalent and thus for simplicity we will use the same symbol for both representations. Next, we investigate the commutant of the representations $U_{g}\otimes\bar{U}_{g}$ and $U_{g}\otimes U_{g}$ in $\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}\cong\mathbb{C}^{n^{2}\times n^{2}}$, i.e. we are interested in the solutions $Z$ of
$(U_{g}\otimes\bar{U}_{g})Z=Z(U_{g}\otimes\bar{U}_{g})\quad\text{for all}\quad g\in G$ (12)
@@
respectively. Eq. (12) and (13) are equivalent to $Ad_{U_{g}\otimes\bar{U}_{g}}(Z)=Z$ and $Ad_{U_{g}\otimes U_{g}}(Z)=Z$, respectively. Finally, a tedious but straightforward computation shows that the partial transposed operator $\Phi:\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}\to\mathbb{C}^{n\times n}\otimes\mathbb{C}^{n\times n}$, $A\otimes B\mapsto A\otimes B^{T}$ yield an intertwining map for $Ad_{U_{g}\otimes\bar{U}_{g}}$ and $Ad_{U_{g}\otimes U_{g}}$, i.e.
$Ad_{U_{g}\otimes\bar{U}_{g}}\circ\Phi=\Phi\circ Ad_{U_{g}\otimes U_{g}}\,.$
This implies $\Phi$ maps the commutator of $\tilde{\pi}$ to the commutator of $\pi\otimes\pi$ and therefore the dimensions of commutators (and consequently the number of irreducible subspaces) coincide. Thus $\mathrm{Sym}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ and $\mathrm{Alt}(\mathbb{C}^{n}\otimes\mathbb{C}^{n})$ are the only irreducible subspaces of $\pi\otimes\pi$ if and only if $\mathbb{C}\,I_{n}$ and $\mathrm{sl}_{0}(n)$ are the only irreducible subspaces to $\tilde{\pi}$ (or, equivalently, $i\mathfrak{su}(n)$ is the only irreducible subspace of $\hat{\pi}$). ∎
@@
Note that the above result does not mean that the tensor square representation $\pi\otimes\pi$ and $\tilde{\pi}$ are equivalent and, in fact, they are not—as easy examples demonstrate. A similar result for Lie-algebra representations was elaborated on in [65] as follow-up on [64] and in particular on [17].
Lemma 5.5.
@@
By Lemma 5.3 we know that the $\hat{\pi}$-invariant subspace spanned by $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ has to coincide with $i\mathfrak{su}(n)$ and thus $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ contains a basis of $i\mathfrak{su}(n)$. Moreover, by the defining property (b) it is straightforward to see that every $1$- and therefore also every $2$-design has to act irreducibly on $\mathbb{C}^{n}$. Thus we conclude
@@
because the left hand side of the above equation belongs to the commutant of $\pi$ and has trace zero. Now if there is some $g_{0}\in G$ such that $\{U_{g}H_{s}U_{g}^{*}:g\in G\,,g\neq g_{0}\}$ does not span $i\mathfrak{su}(n)$, then $U_{g_{0}}H_{s}U_{g_{0}}^{*}$ does not lie in the span of the others, and hence $\sum_{g\in G}U_{g}H_{s}U_{g}^{*}\neq 0$, which contradicts the above. ∎
@@ 6 Conclusion, Discussion, and Outlook
We have analyzed the convergence properties of a recently introduced Haar-randomly projected gradient descent algorithm for cost functions taking the form of a smooth Morse–Bott function (with compact sublevel sets) on a Riemannian manifold. For making it efficient in quantum optimizations, one can approximate the Haar-random projections via unitary 2-designs. For both scenarios we have proven that (i) the respective algorithm almost surely escapes saddle points (Lem. 3.11) and (ii) it almost surely converges to a local minimum (Thm. 3.14). Moreover we have studied the time required by the algorithm to pass a saddle point in a simple two-dimensional setting (Sec. 4.1). Note that unlike adiabatic ground-state preparation strategies (that rest on first preparing some desired initial state to find the ground state of some target Hamiltonian), our randomized Riemannian gradient flow algorithm does not require knowledge of the initial state, as the algorithm converges for almost all initial points. Moreover, as in the quantum setting for ground state problems the critical points just comprise saddles and global extrema, here our result implies almost sure convergence to the global minimum.
@@
The algorithmic steps analyzed here are entirely modular. In view of future applications the (approximate) random projections w.r.t. the Haar measure may well be replaced by selections from other problem-adapted measures without sacrificing the convergence properties. In particular, one may wish to select the measures such that the projections of the gradient do not subside in numerical noise prematurely and thus circumvent the notorious barren-plateaux problem. Moreover, one may consider random projections into subspaces by techniques of compressed sensing and shadow tomography to better approximate the full gradient flow. Likewise, approximations with tensor-network methods (such as, e.g., [60, 61]) may be envisaged as long as one remains on a Riemannian manifold. Needless to say the techniques presented here lend themselves to be taken over to higher-order quasi Newton methods (like L-BFGS) also on quantum computers.
Thus we anticipate that the convergence guaranteed for (randomized) Riemannian gradient flows w.r.t. Morse–Bott type smooth cost functions will encourage wide application in and beyond quantum optimization.
@@ Appendix A Some Technical Results
The probability distributions of $u_{N}$ and $u_{N}^{2}$ are important for understanding the behaviour of the algorithm. Fortunately they can be described quite easily using the beta distribution. Recall that the p.d.f of the beta distribution with parameters $a,b$ is given by $f(x)=\mathrm{const}\cdot x^{a-1}(1-x)^{b-1}.$
@@
If $u\sim\mathcal{U}(S^{N-1})$ and $u_{N}$ is the last coordinate, then
$u_{N}\sim\mathcal{B}_{N}:=2\,\mathrm{Beta}\left(\frac{N-1}{2},\frac{N-1}{2}\right)-1,\quad u_{N}^{2}\sim\mathcal{B}_{N}^{2}:=\mathrm{Beta}\left(\frac{1}{2},\frac{N-1}{2}\right)$
@@
$d_{K}(\sqrt{N}\mathcal{B}_{N},\mathcal{N}(0,1))\leq\frac{1}{N}$
@@
The first inequality follows from [51, p. 20] and [50, Thm. 1.2]. It is easy to see that taking the square of both distributions at most doubles the Kolmogorov distance, and this yields the second inequality. Finally we compute
$\Pr\Big{(}u^{2}_{N}\geq\frac{1}{k^{2}N}\Big{)}=2\big{(}1-\Pr\big{(}u_{N}\leq\tfrac{1}{k\sqrt{N}}\big{)}\big{)}\geq 2\big{(}1-\Phi\big{(}\tfrac{1}{k}\big{)}-\frac{1}{N}\big{)},$
@@
Lemma A.3.
Let $M$ be a compact manifold and let $\{v_{j}\}_{j=1}^{k}$ be a set of continuous vector fields on $M$ which span the tangent space at every point. Now let any vector field $v$ on $M$ be given. Then it holds that there is some $\varepsilon>0$ such that
$\min_{x\in M}\max_{i}\langle v_{i},v\rangle^{2}=\varepsilon.$
Proof.
The functions $\langle v_{j},v\rangle^{2}$ defined on $M$ are smooth and non-negative. Hence their maximum over $j$ is still a continuous non-negative function on $M$. Since the $v_{j}$ span each tangent space, this function is even strictly positive, and by compactness of $M$, there is a positive global minimum. ∎

### v2

@@ 5 Major Application:
To continue the mathematical perspective of ‘ground state’ optimization, i.e. of finding the smallest eigenvalue (‘ground state energy’) of a Hermitian operator $A$ (Hamiltonian) that describes the energy of the system, we leave the general notation $f:M\to\mathbb{R}$ above and specialize to the cost function over the unitary group (or orbit) $J:\mathrm{U}(d)\to\mathbb{R}$ we aim to minimize. For finding the smallest eigenvalue or, more generally, the smallest expectation value of $A$ with respect to the initial state $\rho$, the cost function is given by
$J(U):=\operatorname{tr}(AU\rho U^{*})=:\langle A,\operatorname{Ad}_{U}(\rho)\rangle,$ (11)
@@
where $\rho$ is a density matrix representing the initial state of the system and $U$ is a unitary transformation that describes a quantum circuit (writing $U^{*}$ for the complex conjugate transpose).
@@
By choosing an appropriate eigenbasis of $A$, we may assume that $A$ is diagonal with eigenvalues in non-increasing order. The function we want to optimize is of the form of Eqn. (11). Many properties of this function can be found in [26] in the more general setting of semisimple Lie groups. We recall the relevant properties here, adapted to our setting:
@@
Assume that $\rho$ and $A$ are diagonal  [^12] This amounts to a shift in the function $J$ which results from an appropriate choice of $U:=U_{0}^{*}V_{0}$ and the identity $\operatorname{tr}(A(U_{0}^{*}V_{0})\rho(U_{0}^{*}V_{0})^{*})=\operatorname{tr}(U_{0}AU_{0}^{*}V_{0}\rho V_{0}^{*})$.. Then the following hold.
@@
    $U$ is a critical point of $f$ if and only if $[A,\operatorname{Ad}_{U}(\rho)]=0$. Hence, the critical set of $J$ is equal to
    $\mathrm{SU}(d)_{A}\,\mathcal{S}_{d}\,\mathrm{SU}(d)_{\rho},$
    where $\mathcal{S}_{d}$ is the set of $d$ by $d$ permutation matrices and $\mathrm{SU}(d)_{A}$, $\mathrm{SU}(d)_{\rho}$ denote the stabilizer subgroups for the action of conjugation. In particular, it is a disjoint union of finitely many compact connected submanifolds.
@@
    The Hessian at a critical point $U$ is determined by
@@
    and in particular $J$ is Morse–Bott. [^13] Note that the function $J$ is never Morse, since the maximal torus in $\mathrm{SU}(d)$ stabilizes $A$ and $\rho$.
@@
i: See Lem. 1.1 and Props. 1.2 and 1.3 of [26]. ii: See Prop. 1.4 and Cor. 1.5 of [26]. iii: See Rmk. 1.6 of [26]. ∎
A key tool for the experimental implementation of the randomized gradient descent algorithm on quantum hardware are unitary $t$-designs, whose purpose is to efficiently implement random unitaries.
Definition 5.2 (Thm. 3 in [30]).
A unitary representation $\pi:G\to\mathrm{U}(d)$, $g\mapsto\pi(g)=:U_{g}$ of a finite group $G$ is called a unitary $t$-design with $t\in\mathbb{N}_{0}$ if one of the following equivalent conditions is satisfied:
@@
    where $\mathbb{C}_{t}[Z,\bar{Z}]$ denotes the set of all polynomials in $Z:=(z_{ij})_{i,j=1,\dots d}$ and $\bar{Z}:=(\bar{z}_{ij})_{i,j=1,\dots d}$, which are $t$ homogeneous in $Z$ and $\bar{Z}$., i.e. [^14] This is equivalent to $p(\lambda Z,\bar{\lambda}\bar{Z})=|\lambda|^{2t}p(Z,\bar{Z})$. with $p(\lambda Z,\mu\bar{Z})=\lambda^{t}\mu^{t}p(Z,\bar{Z})$.
@@
    For all $H\in\mathbb{C}^{d^{t}\times d^{t}}$ one has the equality
@@
    The tensor square representation $\pi\otimes\pi:G\to\mathrm{U}(d)\otimes\mathrm{U}(d)$, $g\mapsto U_{g}\otimes U_{g}$ acting on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ has exactly two irreducible components, namely $\mathrm{Sym}(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ and $\mathrm{Alt}(\mathbb{C}^{d}\otimes\mathbb{C}^{d}):=\mathbb{C}^{d}\wedge\mathbb{C}^{d}$.
@@
A more general notion of $t$-designs [24, 25] (see also [30]) allows any finite subset of $\mathrm{U}(d)$ which satisfies (a), (b) or, equivalently,(c). However, since almost all known examples are constructed via group representations, we focus here on this restricted approach which is sometimes called group design. Moreover, the reader should note the following to facts: (i) One can assume without loss of generality that $\pi$ is faithful (i.e., one-to-one) because it is straightforward to show that $[g]\mapsto\pi(g)$, $[g]\in G/\operatorname{ker}\pi$ yields a $t$-design whenever $\pi$ is a $t$-design. (ii) Every $t$-design is also $t^{\prime}$-design for $t^{\prime}\leq t$. This follows readily from condition (b) by choosing $H$ of the form $H^{\prime}\otimes I_{n}\otimes\cdots\otimes I_{n}$.
Now we need the following technical lemma:
@@
Let $\pi:G\to\mathrm{U}(d)$ be any unitary representation. Then $\pi$ yields a $2$-design in the sense of Def. 5.2 if and only if $\pi$ acts irreducibly on $i\mathfrak{su}(n)$ via conjugation, i.e. the representation $\hat{\pi}:G\to U\big{(}i\mathfrak{su}(d)\big{)}$, $\pi(g)\,H:=U_{g}HU_{g}^{*}$ is irreducible.
@@
First, $\hat{\pi}$ obviously “extends” to $\tilde{\pi}:G\to U\big{(}\mathbb{C}^{d\times d}\big{)}$, $\tilde{\pi}(g)\,A:=U_{g}AU_{g}^{-1}$. Then a straightforward computation shows that the representations $\tilde{\pi}$ and $g\mapsto U_{g}\otimes\bar{U}_{g}$ are equivalent and thus for simplicity we will use the same symbol for both representations. Next, we investigate the commutant of the representations $U_{g}\otimes\bar{U}_{g}$ and $U_{g}\otimes U_{g}$ in $\mathbb{C}^{d\times d}\otimes\mathbb{C}^{d\times d}\cong\mathbb{C}^{d^{2}\times d^{2}}$, i.e. we are interested in the solutions $Z$ of
$(U_{g}\otimes\bar{U}_{g})Z=Z(U_{g}\otimes\bar{U}_{g})\quad\text{for all}\quad g\in G\,,$ (12)
@@
respectively. Eq. (12) and (13) are equivalent to $\operatorname{Ad}_{U_{g}\otimes\bar{U}_{g}}(Z)=Z$ and $\operatorname{Ad}_{U_{g}\otimes U_{g}}(Z)=Z$, respectively. Finally, a tedious but straightforward computation shows that the partial transposed operator $\Phi:\mathbb{C}^{d\times d}\otimes\mathbb{C}^{d\times d}\to\mathbb{C}^{d\times d}\otimes\mathbb{C}^{d\times d}$, $A\otimes B\mapsto A\otimes B^{T}$ yield an intertwining map for $\operatorname{Ad}_{U_{g}\otimes\bar{U}_{g}}$ and $\operatorname{Ad}_{U_{g}\otimes U_{g}}$, i.e.
$\operatorname{Ad}_{U_{g}\otimes\bar{U}_{g}}\circ\,\Phi=\Phi\circ\operatorname{Ad}_{U_{g}\otimes U_{g}}\,.$
This implies $\Phi$ maps the commutator of $\tilde{\pi}$ to the commutator of $\pi\otimes\pi$ and therefore the dimensions of commutators (and consequently the number of irreducible subspaces) coincide. Thus $\mathrm{Sym}(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ and $\mathrm{Alt}(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ are the only irreducible subspaces of $\pi\otimes\pi$ if and only if $\mathbb{C}\,I_{d}$ and $\mathfrak{sl}_{0}(d)$ are the only irreducible subspaces to $\tilde{\pi}$ (or, equivalently, $i\mathfrak{su}(d)$ is the only irreducible subspace of $\hat{\pi}$). ∎
@@
Note that the above result does not mean that the tensor square representation $\pi\otimes\pi$ and $\tilde{\pi}$ are equivalent and, in fact, they are not — as easy examples demonstrate. A similar result for Lie algebra representations was elaborated on in [75] as follow-up to [74] and in particular to [20].
The following is the main result of this section, showing that unitary $2$-designs satisfy Assumption 3.1. As a consequence, the main convergence result (Thm. 3.17) applies to the problem considered in this section.
Proposition 5.6.
@@
By Lem. 5.4 we know that the $\hat{\pi}$-invariant subspace spanned by $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ has to coincide with $i\mathfrak{su}(d)$ and thus $\{U_{g}H_{s}U_{g}^{*}:g\in G\}$ contains a basis of $i\mathfrak{su}(d)$. Moreover, by the defining property (b) it is straightforward to see that every $1$-design, and therefore also every $2$-design, has to act irreducibly on $\mathbb{C}^{d}$. Thus we conclude
@@
because the left hand side of the above equation belongs to the commutant of $\pi$ and has trace zero. Now if there is some $g_{0}\in G$ such that $\{U_{g}H_{s}U_{g}^{*}:g\in G\,,g\neq g_{0}\}$ does not span $i\mathfrak{su}(d)$, then $U_{g_{0}}H_{s}U_{g_{0}}^{*}$ does not lie in the span of the others, and hence $\sum_{g\in G}U_{g}H_{s}U_{g}^{*}\neq 0$, which contradicts the above. ∎
Corollary 5.7.
The randomized gradient descent algorithm introduced in Sec. 3, applied to the cost function (11) on the unitary group, with randomly chosen directions $\{U_{g}H_{s}U_{g}^{*}\}_{g\in G}$ as in Prop. 5.6, converges almost surely to a global minimum. Moreover, by continuity, this continues to hold if $U_{g}$ is an approximate $2$-design, as long as the approximation is good enough.
On the Orbit
Since the map $J:\mathrm{SU}(d)\to\mathbb{R}$, which depends on the initial state $\rho$, is invariant under the stabilizer subgroup of $\rho$, denoted $\mathrm{SU}(d)_{\rho}$, the map can be factored through the corresponding quotient space. Equivalently, one may define the map on the unitary orbit of the initial state $\rho$. We denote this map by $\bar{J}:\mathrm{SU}(d)\rho\to\mathbb{R}$. Working with this function has some theoretical advantages, such as the reduced dimension of the state space. Moreover, in many cases the function is even Morse:
Lemma 5.8.
If $A$ is non-degenerate, then the function $\bar{J}$ is Morse.
Proof.
See [26, Coro. 3.7, Sec. 4]. ∎
Despite the nice properties of $\bar{J}$, we do not use this formulation, since it is not clear how to efficiently sample random unitaries yielding directions in the tangent space of the orbit.
@@ 6 Conclusion, Discussion, and Outlook
We have analyzed the convergence properties of a recently introduced Haar-randomly projected gradient descent algorithm for cost functions taking the form of a smooth Morse–Bott function (with compact sublevel sets) on a Riemannian manifold. For making it efficient in quantum optimizations, one can approximate the Haar-random projections via unitary 2-designs. For both scenarios we have proven that (i) the respective algorithm almost surely escapes saddle points (Lem. 3.14) and (ii) it almost surely converges to a local minimum (Thm. 3.17). Moreover, we have studied the time required by the algorithm to pass a saddle point in a simple two-dimensional setting (Sec. 4.1). Note that unlike adiabatic ground-state preparation strategies (that rest on first preparing some desired initial state to find the ground state of some target Hamiltonian), our randomized Riemannian gradient flow algorithm does not require knowledge of the initial state, as the algorithm converges for almost all initial points. Furthermore, as in the quantum setting for ground state problems the critical points just comprise saddles and global extrema, here our result implies almost sure convergence to the global minimum.
@@
The algorithmic steps analyzed here are entirely modular. In view of future applications, the (approximate) random projections w.r.t. the Haar measure may well be replaced by selections from other problem-adapted measures without sacrificing the convergence properties. In particular, one may wish to select the measures such that the projections of the gradient do not subside in numerical noise prematurely and thus circumvent the notorious barren plateaux problem. Moreover, one may consider random projections into subspaces by techniques of compressed sensing and shadow tomography to better approximate the full gradient flow. Likewise, approximations with tensor-network methods (such as, e.g., [70, 71]) may be envisaged as long as one remains on a Riemannian manifold. More specifically, see [52] for the connection of trotterized MERA tensor networks with Riemannian optimization (extended to quasi-Newton) of VQAs in view of hybrid quantum computing. Another interesting recent application, which may potentially profit from the random methods described here, is Riemannian quantum circuit optimization based on matrix product operators as suggested by [41].
Needless to say, the randomized gradient techniques presented here in this paper lend themselves to be taken over to higher-order quasi Newton methods (like L-BFGS) as described in [63] also on quantum computers.
Thus we anticipate that the convergence guaranteed for (randomized) Riemannian gradient flows with respect to Morse–Bott type smooth cost functions will encourage wide application in and beyond quantum optimization.
@@ Acknowledgements
Conflicts of Interest
All authors declare that they have no conflicts of interest.
@@ Appendix A Some Technical Results
The probability distributions of $u_{N}$ and $u_{N}^{2}$ are important for understanding the behavior of the algorithm. Fortunately, they can be described quite easily using the beta distribution. Recall that the p.d.f. of the beta distribution with parameters $a,b>0$ is given by $f(x)=\mathrm{const}\cdot x^{a-1}(1-x)^{b-1}$.
@@
If $\mathbf{u}\sim\mathcal{U}(S^{N-1})$ and $u_{N}$ is the last coordinate, then
$u_{N}\sim\mathcal{B}_{N}:=2\,\mathrm{Beta}\left(\frac{N-1}{2},\frac{N-1}{2}\right)-1,\quad u_{N}^{2}\sim\mathcal{B}_{N}^{2}:=\mathrm{Beta}\left(\frac{1}{2},\frac{N-1}{2}\right),$
@@
$d_{K}(\sqrt{N}\mathcal{B}_{N},\mathcal{N}(0,1))\leq\frac{1}{N}\,,$
@@
The first inequality follows from [60, p. 20] and [59, Thm. 1.2]. It is easy to see that taking the square of both distributions at most doubles the Kolmogorov distance, and this yields the second inequality. Finally we compute
$\Pr\big{(}u^{2}_{N}\geq\tfrac{1}{k^{2}N}\big{)}=2\big{(}1-\Pr\big{(}u_{N}\leq\tfrac{1}{k\sqrt{N}}\big{)}\big{)}\geq 2\big{(}1-\Phi\big{(}\tfrac{1}{k}\big{)}-\tfrac{1}{N}\big{)},$